# Shaken, Not Stirred: A Novel Dataset for Visual Understanding of Glasses in Human-Robot Bartending Tasks

### v2

@@ III-E Glass Base Points
To improve the robotic arm guidance without heavy modification of the detector architecture or lifting to 3D [49], we calculate 2D glass base positions separately. Given these locations and the known heights of the glasses from the classification, it enables robot pouring without exact 3D positions. We provide annotations for them in the dataset. By calculating the average position of each 3D-printed cap, we obtain point $\mathbf{p}=(x,y,z)^{\top}$ for each glass and collect them into a set $\mathbf{C}$. Taking the normal vector of the table $\mathbf{n}=(a,b,c)^{\top}$ and its normalized version $\mathbf{\hat{n}}$, the perpendicular distance to the table for each $\mathbf{p}\in\mathbf{C}$ is given by
@@
$H(x,y)=\sum_{i}s_{i}G_{i}(x,y)$ (5)
The 2D base points are extracted by identifying local maxima in the heatmap within each detected glass bounding box. See Fig. 2(d) for an example network output, with heatmap overlayed on glass detections. These points can also be calculated with the approach described in Section III-G. Full 3D localization using the depth maps provided in that dataset is proposed for future work.
@@ 1. Pouring Wine:
- Example: $<$pour_wine(wine glass)$>$.
@@ 2. Pouring Beer:
- Example: $<$pour_beer(beer glass)$>$.
- Example: $<$pour_beer(high beer glass)$>$.
@@ III-F LLM Integration: Bartender Task
The NICOL platform supports an LLM integration for embodied agents in real-world robotic tasks [51]. We utilize the grounded LLM to implement a robot bartender scenario, leveraging the LLM’s reasoning to connect the robot’s sensory perception (identifying glass objects) and robot’s physical actions (precise arm manipulation for pouring). Our task facilitates natural interaction with users since the bartender scenario inherently involves social dynamics like recognizing drinkware, interpreting user requests, and responding appropriately. Although our implementation focuses on verbal communication, the task can be extended to incorporate context-driven and multimodal cues, including nonverbal communication. The LLM prompt is in Prompt III-E.
@@ III-G Robot Integration: Beverage Pouring Action
The Bézier curve-based motion planner described in our previous work [52] is utilized to design a pouring motion for the robot bartender task. The action picks up an already open bottle from a fixed position, moves it towards the opening of the glass, pours particles into it, and returns the bottle to its origin. Careful and precise pouring motions are required to prevent spilling particles and to avoid damaging the glasses [^1] No glass was harmed in any of our experiments.. The RealSense camera attached to the head of NICOL is used for this application, but it is also possible to use the left or right eye cameras. We use 3D-printed particles to imitate liquid as it prevents damage to the electrical hardware.
@@
$\begin{split}p_{x}=\epsilon\cdot p_{x}^{min}+(\epsilon\cdot\tau)p_{x}^{max}\\ p_{y}=\gamma\cdot p_{y}^{min}+(\gamma\cdot\tau)p_{y}^{max}\end{split}$ (6)
In the prior equation, suitable offsets for the smallest glass class $p_{x}^{min}$ and $p_{y}^{min}$ are adjusted with the linear scaling factor for the x-axis position $\epsilon$ and the y-axis position $\gamma$. In addition, suitable offsets for the highest glass class $p_{x}^{max}$ and $p_{y}^{max}$ are added by adjusting it with the linear scaling factor for the glass height $\tau$. As the height offset has shown to be influenced by the x- and y-position of the glass in our preliminary experiments, it is multiplied by the scaling factor for the corresponding axis $\epsilon$ or $\gamma$. The pouring offsets for the smallest and highest glass classes $p_{x}^{min}$, $p_{y}^{min}$, $p_{x}^{max}$, and $p_{y}^{max}$ can be determined by placing the smallest and highest glass at the maximum x- and y-position of the pouring workspace and manually tuning for the suitable offsets. The primitive scales well for the intermediate glass classes. The normalized scaling factors $\epsilon$, $\gamma$, and $\tau$ are calculated according to Eq. 7:
$\begin{split}\epsilon=&\frac{x_{n}-x_{min}}{x_{max}-x_{min}},\hskip 5.0pt\gamma=\frac{|y_{n}|-y_{min}}{y_{max}-y_{min}}\\ &\hskip 20.0pt\tau=\frac{h_{n}-h_{min}}{h_{max}-h_{min}}\\ \end{split}$ (7)
@@ IV Experiments and Evaluation
We conduct two experiments to evaluate our system. First, we evaluate the performance of our fine-tuned RTMDet glass classifier using our proposed dataset and compare it against state-of-the-art off-the-shelf open-vocabulary object detectors. In the second experiment, we integrate our glass classifier with an LLM-powered embodied agent using the NICOL robot and perform an end-to-end system evaluation in a bartender-like Human-Robot scenario. In the next subsections, we show the evaluation metrics and results of each experiment.
@@ IV-A Transparent Object Detection
We evaluate the detector on the test data using the two standard COCO evaluation metrics Average Precision and Average Recall at IoU from $0.5$ to $0.95$. Using data from the eye cameras in addition to the captures from the head-mounted RealSense during training slightly increases the performance. The values are calculated for two cases. In the first case (highlighted in Table I), we use the class names described in Sec. III-B as prompts to the open-vocabulary detectors. The shown values of AP and AR are the mean across all glass types (sometimes referred to as mAP and mAR, respectively). It should be mentioned that existing detectors are unaware of our classification of glass types. There is an ambiguity even for humans, such as the difference between a whisky and a water glass. Therefore, in the second case, we treat all instances as a general drink glass class. In preliminary experiments, we have also tried other prompts, such as transparent drink container, drinking glass, or drinkware with worse results. The testing split, as the rest of our data, does not contain any other objects classifiable as a drink glass by a human. We have also experimented with image-guided prompting with a template glass image. In such a configuration, the detectors preferred to look for other containers in a similar pose (such as from a top-right view), rather than focusing on the material and type. This caused an even higher false positive rate, with detections of objects like ceramic cups in a pose similar to the template.
The higher AP of our retrained model, even in the general drink glass setting, quantitatively proves our hypothesis that existing open-vocabulary detectors do not understand the concept of a drink container made from a glass material. In their latent space, it seems to be confused with containers made from arbitrary different materials, such as a metal cans. They also cannot distinguish between transparent containers like bottles and drinking glasses, whereas for humans they serve different purposes. Going one additional step towards glass type classification makes existing detectors basically unusable for our scenario. In terms of AR, existing models perform more appropriately. They do not have many false negatives, generally selecting everything that looks like a rounded container.
@@ IV-B Robot Integration: Bartender Task
In this experiment, we use the LLM integration (see Sec. III-F) and beverage pouring motion (see Sec. III-G) to evaluate the performance of our robot bartender task, deployed on the physical robot hardware. We evaluate our system in an end-to-end manner, i.e., a user (an experimenter), interacts with the system in a way that resembles a real-world scenario. Our experimental setup consists of NICOL with six glass objects placed on the table – a shot glass, wine glass, whiskey glass, water glass, beer glass, and high beer glass. Additionally, two plastic bottles – a wine bottle and beer bottle – are placed on each side of the table and filled with artificial liquid made from 3D-printed round-shaped particles with distinct colors. The pouring workspace has a width of 55 cm and a depth of 35 cm. The positions of the bottles were swapped halfway through the experiment for a fair comparison of the pouring by each arm. The user stands across the table and issues verbal commands through an external microphone, processed via a Whisper integration.
@@
Each interaction consists of the following: the user orders wine, then a beer, followed by another beer request. We select these interactions to assess the system’s ability to adapt to different beverage types, resolve ambiguities in glassware selection, and ensure a logical action sequence. This can be noted as a three-phase process: 1)  User Order Recognition: The system processes the user’s order and determines the appropriate drink and glassware, through action-object mapping or user preference. For example: if the user says “I’d like a wine”, the robot must infer that wine should be poured into a wine glass, 2) Reasoning & Decision Making: Based on user input, the robot reasons about the correct bottle and glass object. Any ambiguity is clarified through user feedback. For example: if the user says “Can I get a beer for my friend?”, the robot must decide the correct glassware based on context and prior interactions (which glass is already filled), 3) Action Execution: The robot performs a pouring action correctly to serve the requested drink. We use OpenAI GPT‑4o mini and reset the chat after each interaction.
@@
We conducted 35 trials, meaning that 105 pouring actions were executed in total. We use the Success rate of the end-to-end pouring action as a metric to evaluate our experiment. However, we report the average error rate in the following categories: Spill referring to slight spilling of particles outside the intended glass, LLM representing LLM reasoning inaccuracies, and Detection, i.e., errors in the glass detector (cf. Fig. 3). Each pouring motion is considered successful if the robot correctly pours the target drink into the target glass without any particles spilled. Our system achieves a success rate of 81% in the pouring action across the two robot arms, while only 4.8% of the cases exhibited minor spills of particles outside the designated glass, demonstrating the suitability of our approach for precise pouring in a bartender task. Our LLM integration showed high resilience in generating correct robot actions across the different use cases: action-object mapping, ambiguity resolution, and logical action sequencing. In 10.5% of the cases, user correction was needed due to reasoning inconsistencies, such as suggesting the wrong glass for a specific drink. However, only one instance was recorded, where the LLM generated an unintended action, mistakenly handing over the object instead of pouring into it, leading to a failure case. Only 4.8% of the cases resulted in failures due to occasional glass detector errors, such as failing to detect the beer glass.
@@
V Discussion
As can be seen in Table I, our new benchmark dataset helps to push the current state-of-the-art detector boundaries. Zero-shot models like YOLO-World[7], OWL-Vit [8], and GroundingDINO[6] are widely used for opaque models, but their performance deteriorates on concepts of transparency, glass material and drink-purposed containers. On the other hand, our processing pipeline allows us to train a lightweight model from RTMDet [40] family, achieving a high precision with a smaller rate of false positives. This proves the fact that reliable recognition of glass objects is far from achieved with general models. The glass detection system has shown high performance when integrated with the physical robot, achieving an overall success rate of 81% in a bartending scenario. While the results show that there is still room for improvements in our object detection and robotic action modules, the biggest error source originates from the LLM utilized for high-level planning, making more than 50% of the failure trials. Unfortunately, the glass base keypoints described in Sec. III-E oscillated too much to be applicable in the bartending agent in preliminary experiments. However, 3D planning approaches seem to be a promising area for future research. Our system already shows high performance by planning the motion based on the 2D image perception of the scene. A 3D perception approach, e.g., by utilizing the chalk samples delivered with our dataset for depth estimation, would thus still be capable of improving it.
@@ VI CONCLUSION
This paper fills a critical gap in glass object perception by providing a novel real-world dataset. The proposed auto-labeling method reduces the human effort for data aggregation in similar tasks to an absolute minimum. We contribute three distinct repositories under an open-source license: the glass dataset, the glass detector model and experiment code, and the NICOL camera model. The quality of our dataset and camera model allows for very precise projections between different camera perspectives. Our glass detector outperforms SOTA open-vocabulary approaches and reaches a baseline success rate of 81% when integrated with our embodied bartending agent. Future work will include expanding the dataset with more diverse scenes and improving robotic manipulation strategies through enhanced LLM integration. The most interesting next step for the action module is a proprioceptive 3D planning approach that can dynamically adapt to spatial changes, like a user moving a glass while pouring, fostering seamless Human-Robot collaboration in real-world tasks.

### v3

@@ III-E Glass Base Points
To improve the robotic arm guidance without heavy modification of the detector architecture or lifting to 3D [51], we calculate 2D glass base positions separately. Given these locations and the known heights of the glasses from the classification, it enables robot pouring without exact 3D positions. We provide annotations for them in the dataset. By calculating the average position of each 3D-printed cap, we obtain point $\mathbf{p}=(x,y,z)^{\top}$ for each glass and collect them into a set $\mathbf{C}$. Taking the normal vector of the table $\mathbf{n}=(a,b,c)^{\top}$ and its normalized version $\mathbf{\hat{n}}$, the perpendicular distance to the table for each $\mathbf{p}\in\mathbf{C}$ is given by:
@@
$H(x,y)=\sum_{i}s_{i}G_{i}(x,y).$ (5)
2D base points are extracted by identifying local maxima in the heatmap within each detected glass bounding box (see Fig. 2(d) for an example with heatmap overlaid on detected glasses). 2D points can also be calculated with the approach in Section III-G. Full 3D localization using the depth maps provided in that dataset is proposed for future work.
@@ 1. Pouring Wine:
- Example: <<pour_wine(wine glass)>>.
@@ 2. Pouring Beer:
- Example: <<pour_beer(beer glass)>>.
- Example: <<pour_beer(high beer glass)>>.
@@ III-F LLM Integration: Bartender Task
The NICOL platform supports an LLM integration for embodied agents in real-world robotic tasks [53]. We utilize the grounded LLM to implement a robot bartender scenario, leveraging the LLM’s reasoning to connect the robot’s sensory perception (identifying glass objects) and robot’s physical actions (precise arm manipulation for pouring). Our task facilitates natural interaction with users since the bartender scenario inherently involves social dynamics like recognizing drinkware, interpreting user requests, and responding appropriately. Moreover, bartending tasks have proven to be engaging testbeds for social HRI, as shown in systems like BRILLO [54]. Although our implementation focuses on verbal communication, the task can be extended to incorporate context-driven and multimodal cues, including nonverbal communication. The LLM prompt is in Prompt III-E.
@@ III-G Robot Integration: Beverage Pouring Action
The Bézier curve-based motion planner described in our previous work [55] is utilized to design a pouring motion for the robot bartender task. The action picks up an already open bottle from a fixed position, moves it towards the opening of the glass, pours particles into it, and returns the bottle to its origin. Careful and precise pouring motions are required to prevent spilling particles and to avoid damaging the glasses [^1] No glass was harmed in any of our experiments.. The RealSense camera attached to the head of NICOL is used for this application, but it is also possible to use the left or right eye cameras. We use 3D-printed particles to imitate liquid, as it prevents damage to the electrical hardware.
@@
$\begin{split}p_{x}=\epsilon\cdot p_{x}^{min}+(\epsilon\cdot\tau)p_{x}^{max}\\ p_{y}=\gamma\cdot p_{y}^{min}+(\gamma\cdot\tau)p_{y}^{max}.\end{split}$ (6)
In the prior equation, suitable offsets for the smallest glass class $p_{x}^{min}$ and $p_{y}^{min}$ are adjusted with the linear scaling factor for the x-axis position $\epsilon$ and the y-axis position $\gamma$. In addition, suitable offsets for the highest glass class $p_{x}^{max}$ and $p_{y}^{max}$ are added by adjusting them with the linear scaling factor for the glass height $\tau$. As the height offset has shown to be influenced by the x- and y-position of the glass in our preliminary experiments, it is multiplied by the scaling factor for the corresponding axis $\epsilon$ or $\gamma$. The pouring offsets for the smallest and highest glass classes $p_{x}^{min}$, $p_{y}^{min}$, $p_{x}^{max}$, and $p_{y}^{max}$ are determined by placing the smallest and highest glass at the maximum x- and y-position of the pouring workspace and manually tuning the offsets. The primitive scales well to intermediate glass classes. The normalized scaling factors $\epsilon$, $\gamma$, and $\tau$ are calculated as in Eq. 7:
[…]
@@ IV Experiments and Evaluation
We conduct two experiments to evaluate our system. First, we evaluate the performance of our fine-tuned RTMDet glass classifier using our proposed dataset and compare it against state-of-the-art off-the-shelf open-vocabulary object detectors. Additionally, we verify our claim regarding the insufficiency of synthetic data caused by the challenging glass material. Using 3D models of glasses provided within our dataset, we train the network on data rendered by the SuperCaustics framework [25] and include its performance in the evaluation. In the second experiment, we integrate our glass classifier with an LLM-powered embodied agent using NICOL and perform an end-to-end system evaluation in a bartender-like Human-Robot scenario. Next, we show the evaluation metrics and results of each experiment.
@@ IV-A Transparent Object Detection
We evaluate the detector on the test data using the two standard COCO evaluation metrics Average Precision and Average Recall at IoU from $0.5$ to $0.95$. Using data from the eye cameras in addition to the captures from the head-mounted RealSense during training slightly increases the performance. The values are calculated for two cases. In the first case (highlighted in Table I), we use the class names described in Sec. III-B as prompts to the open-vocabulary detectors. The shown values of AP and AR are the mean across all glass types (sometimes referred to as mAP and mAR, respectively). It should be mentioned that existing detectors are unaware of our classification of glass types. There is an ambiguity even for humans, such as the difference between a whiskey and a water glass. Therefore, in the second case, we treat all instances as a general drink glass class. In preliminary experiments, we have also tried other prompts, such as transparent drink container, drinking glass, or drinkware, with worse results. The testing split, as the rest of our data, does not contain any other objects classifiable as a drink glass by a human. We have also experimented with image-guided prompting with a template glass image. In such a configuration, the detectors preferred to look for other containers in a similar pose (such as from a top-right view), rather than focusing on the material and type. This caused an even higher false positive rate, with detections of objects like ceramic cups in a pose similar to the template.
The higher AP of our pretrained model, even in the general drink glass setting, supports our hypothesis that existing open-vocabulary detectors do not grasp the concept of a drink container made from glass. In their latent space, it seems confused with containers made from other materials like metal cans. They also cannot distinguish between transparent containers like bottles and drinking glasses, which clearly serve different functions. Taking a further step towards glass type classification makes existing detectors unusable for our scenario. In terms of AR, existing models perform more appropriately. They do not have many false negatives, generally selecting everything that looks like a rounded container.
@@ IV-B Robot Integration: Bartender Task
In this experiment, we use the LLM integration (see Sec. III-F) and beverage pouring motion (see Sec. III-G) to evaluate the performance of our robot bartender task, deployed on the physical robot hardware. We evaluate our system in an end-to-end manner, i.e., a user (an experimenter) interacts with the system in a way that resembles a real-world scenario. Our experimental setup consists of NICOL with six glass objects placed on the table – a shot glass, wine glass, whiskey glass, water glass, beer glass, and high beer glass. Additionally, two plastic bottles – a wine bottle and beer bottle – are placed on each side of the table and filled with artificial liquid made from 3D-printed round-shaped particles with distinct colors. The pouring workspace has a width of 55 cm and a depth of 35 cm. The positions of the bottles were swapped halfway through the experiment for a fair comparison of the pouring by each arm. The user stands across the table and issues verbal commands through an external microphone, processed via a Whisper integration.
@@
Each interaction consists of the following: the user orders wine, then a beer, followed by another beer request. We select these interactions to assess the system’s ability to adapt to different beverage types, resolve ambiguities in glassware selection, and ensure a logical action sequence. This follows as a three-phase process: 1)  User Order Recognition: The system processes the user’s order and determines the appropriate drink and glassware, through action-object mapping or user preference. For example: if the user says “I’d like a glass of wine”, the robot must infer that wine should be poured into a wine glass, 2) Reasoning & Decision Making: Based on user input, the robot reasons about the correct bottle and glass object. Any ambiguity is clarified through user feedback. For example: if the user says “Can I get a beer for my friend?”, the robot must decide the correct glassware based on context and prior interactions (which glass is already filled), 3) Action Execution: The robot performs a pouring action correctly to serve the request. We use OpenAI GPT‑4o mini and reset the chat after each interaction.
@@
We conducted 35 trials, meaning that 105 pouring actions were executed in total. We use the Success rate of the end-to-end pouring action as a metric to evaluate our experiment. However, we report the average error rate in the following categories: Spill referring to slight spilling of particles outside the intended glass, LLM representing LLM reasoning inaccuracies, and Detection, i.e., errors in the glass detector (cf. Fig. 3). Each pouring motion is considered successful if the robot correctly pours the target drink into the target glass without any particles spilled. Our system achieves a success rate of 81% in the pouring action across the two robot arms, while only 3.8% of the cases exhibited minor spills of particles outside the designated glass, demonstrating the suitability of our approach for precise pouring in a bartender task. Our LLM integration showed high resilience in generating correct robot actions across the different use cases: action-object mapping, ambiguity resolution, and logical action sequencing. In 10.5% of the cases, user correction was needed due to reasoning inconsistencies, such as suggesting the wrong glass for a specific drink. However, only one instance was recorded, where the LLM generated an unintended action, mistakenly handing over the object instead of pouring into it, leading to a failure case. Only 4.8% of the cases resulted in failures due to occasional glass detector errors, such as failing to detect the beer glass.
V Discussion
As can be seen in Table I, our new benchmark dataset helps to push the current state-of-the-art detector boundaries. Zero-shot models like YOLO-World[8], OWL-Vit [9], and GroundingDINO[7] are widely used for opaque models, but their performance deteriorates on concepts of transparency, glass material, and drink-purposed containers. On the other hand, our processing pipeline allows us to train a lightweight model from the RTMDet [42] family, achieving a high precision with a smaller rate of false positives. This proves the fact that reliable recognition of glass objects is far from achieved with general models. The glass detection system has shown high performance when integrated with the physical robot, achieving an overall success rate of 81% in a bartending scenario. While the results show that there is still room for improvement in our object detection and robotic action modules, the biggest error source originates from the LLM utilized for high-level planning, making more than 50% of the failure trials. Unfortunately, the glass base keypoints described in Sec. III-E oscillated too much to be applicable in the bartending agent in preliminary experiments. However, 3D planning approaches seem to be a promising area for future research. Our system already shows high performance by planning the motion based on the 2D image perception of the scene. A 3D perception approach, e.g., by utilizing the chalk samples delivered with our dataset for depth estimation, would thus still be capable of improving it.
@@ VI CONCLUSION
This paper fills a critical gap in glass object perception by providing a novel real-world dataset. The proposed auto-labeling method reduces the human effort for data aggregation in similar tasks to an absolute minimum. We contribute three distinct open-source repositories: the glass dataset, the glass detector model and experiment code, and the NICOL camera model. The quality of our dataset and camera model allows for very precise projections between different camera perspectives. Our glass detector outperforms SOTA open-vocabulary approaches and reaches a baseline success rate of 81% when integrated with our embodied bartending agent. Future work will include expanding the dataset with more diverse scenes and improving robotic manipulation strategies through enhanced LLM integration. The most interesting next step for the action module is a proprioceptive 3D planning approach that can dynamically adapt to spatial changes, like a user moving a glass while pouring, fostering seamless Human-Robot collaboration in real-world tasks.